### v1

# CLIP-UP: A Simple and Efficient Mixture-of-Experts CLIP Training Recipe with Sparse Upcycling

Abstract
Mixture-of-Experts (MoE) models are crucial for scaling model capacity while controlling inference costs. While integrating MoE into multimodal models like CLIP improves performance, training these models is notoriously challenging and expensive. We propose CLIP-Upcycling ( CLIP-UP), an efficient alternative training strategy that converts a pre-trained dense CLIP model into a sparse MoE architecture. Through extensive experimentation with various settings and auxiliary losses, we demonstrate that CLIP-UP significantly reduces training complexity and cost. Remarkably, our sparse CLIP B/16 model, trained with CLIP-UP, outperforms its dense counterpart by 7.2% and 6.6% on COCO and Flickr30k text-to-image Recall@1 benchmarks respectively. It even surpasses the larger CLIP L/14 model on this task while using only 30% of the inference FLOPs. We further demonstrate the generalizability of our training recipe across different scales, establishing sparse upcycling as a practical and scalable approach for building efficient, high-performance CLIP models.
Machine Learning, ICML
Apple
1 Introduction
CLIP (Radford et al., 2021; Jia et al., 2021) has proven to be a transformative model across a wide range of domains, including image classification, multimodal retrieval, and AI-driven multimodality content generation (Zhou et al., 2022; Rao et al., 2022; Gan et al., 2022; Ramesh et al., 2021; Liu et al., 2023). However, as demands on CLIP grow across these domains, scaling up the model becomes increasingly important to achieve better performance. Recent efforts to scale up CLIP have primarily focused on increasing the size of dense models (Cherti et al., 2023), which, while effective, result in significant computational overhead and high inference costs.
Figure 1: Our proposed Mixture-of-Experts (MoE) CLIP pre-training recipe. We identify several key ingredients for efficient and effective training including shared vs. separated image-text backbones, training from scratch vs. sparse upcycling from a dense model, use of auxiliary losses, and etc. A detailed analysis is provided in Section 4.1. We conclude a simple strategy that outperforms previous practice (Mustafa et al., 2022) with even less overall training cost.
Figure 2: CLIP-UP overview (Left) and sparse upcycling initialization (Right). To upcycle from a pre-trained dense checkpoint (e.g., single expert), the MLP layers in some transformer blocks are replaced with MoE layers. During initialization, the MoE MLP parameters are copied from the original dense checkpoint. The MoE router is randomly initialized.
Another promising approach to scaling CLIP is through sparse modeling, particularly by introducing Mixture-of-Expert (MoE) layers (Mustafa et al., 2022). MoE is a technique that activates only a subset of the model’s experts (parameters) per input, which reduces inference cost (Shazeer et al., 2017) compared to the dense model with the same number of parameters. MoE models have demonstrated great success in various tasks, including multimodal tasks (Fedus et al., 2022; Lepikhin et al., 2020). Despite these advances, there remains key problems in prohibitively high training cost of CLIP like models from scratch and the need of complicated auxiliary losses to stabilize the MoE model training. Mustafa et al. (2022) investigated training the LIMOE model – an MoE CLIP model with a shared backbone for text and image encoders – from scratch. While it significantly outperforms the dense model, LIMOE demands substantial computational resources and introduces newly defined auxiliary losses for enhanced robustness. For instance, LIMOE with CLIP B/16 requires 1.35 times the training FLOPs of its dense counterpart.
In this work, we explore an alternative training strategy: upcycling a pre-trained dense CLIP model to a sparse MoE architecture. Sparse upcycling (Komatsuzaki et al., 2023) leverages a well-trained dense model to initialize a sparse model with MoE layers. Figure 2 illustrates the process in detail. Our experiments start with the CLIP-B/16 architecture from LIMOE, initially trained from scratch with a shared backbone. We then further evaluate the impact of LIMOE auxiliary loss, shared versus separated backbones, and training from scratch versus sparse upcycling. As shown in Figure 1, sparse upcycling with a separated backbone yields the best quality metrics while significantly reducing training costs, lowering ZFLOPS from 4.2 to 3.7 compared to training from scratch. Incorporating LIMOE’s local and global entropy loss  (Mustafa et al., 2022) substantially improves the performance of shared backbone trained from scratch but still lags behind other configurations. Section 4.1 provides a detailed analysis of training strategies and the impact of LIMOE auxiliary loss under different setups.
In particular, we propose CLIP-UP, a sparse upcycling approach (Komatsuzaki et al., 2023) for CLIP models with MoE. As illustrated in Figure 2, this method upgrades a pretrained CLIP into a sparsely activated MoE model with relatively low additional training cost. By leveraging the pretrained dense model, CLIP-UP benefits from a warm start, significantly improving training efficiency. We demonstrate that CLIP-UP outperforms both continued training of the original dense model and training a sparse model from scratch across different scales.
In summary, our contributions are as follows:
  1. 1.
    We introduce CLIP-UP, a simple yet effective training recipe for Mixture-of-Experts (MoE) CLIP models leverages sparse upcycling from pre-trained weights, eliminating the need for complex auxiliary losses. CLIP-UP outperforms existing MoE CLIP training methods, regardless of whether the backbone is shared or separated.
  2. 2.
    CLIP-UP achieves substantial performance improvements across various Text-Image retrieval benchmarks compared to the corresponding dense models. Notably, CLIP-UP with a B/16 backbone surpasses dense CLIP by 7.2% and 5.5% on recall@1 of COCO (Lin et al., 2014) and Flickr30K (Plummer et al., 2017) text-to-image retrieval tasks respectively.
  3. 3.
    We show CLIP-UP is scalable to different model sizes (from B/32 to L/14) and provide a comprehensive analysis of key factors and challenges, providing valuable insights into the design decisions.
2 Related Work
Mixture-of-Experts. Sparsely-activated MoE models have emerged as a powerful approach for scaling model architectures while maintaining computational efficiency (Shazeer et al., 2017; Fedus et al., 2022). By activating only a subset of experts during inference, MoE models significantly reduce the required computation compared to dense models, without sacrificing performance (Jiang et al., 2024; Dai et al., 2024; Xue et al., 2024). However, the additional expert layers and the complexity of routing mechanisms required to select active experts during training make MoE models more resource-intensive to train from scratch.
Early pioneer work like LIMOE (Mustafa et al., 2022) applies the MoE framework to CLIP models, using a shared structure for both text and image modalities. This design enables efficient learning of shared representations, leading to significant performance improvements across various multimodal tasks and outperforming compute-matched dense models.
Despite its advantages, LIMOE shares a common limitation with other MoE models: the need for substantial computational resources during training (Du et al., 2024). As shown in Figure 1, training with a shared backbone from scratch yields suboptimal results compared to alternative backbone selections and training strategies. While LIMOE auxiliary loss offers some improvement, the overall performance remains inferior, suggesting that the methodology may not be the most effective approach.
Sparse upcycling. Sparse upcycling (Komatsuzaki et al., 2023) is a technique that effectively reduce training cost of sparse models by reusing existing dense model checkpoints. By leveraging the prior knowledge from a well-trained dense model, the sparse version benefits from accelerated convergence and reduced computational overhead (He et al., 2024).
The current applications of sparse upcycling have been primarily focused on single-modality models (Komatsuzaki et al., 2023; He et al., 2024). The effectiveness of this approach in the context of multimodal models remains largely unexplored. The concurrent work CLIP-MoE (Zhang et al., 2024) upcycles MoE layers by initializing them with fine-tuned MLP layers via cluster-and-contrast learning. Each expert requires an additional training stage. Scaling this method is challenging, as adding more experts increases complexity and necessitates additional training stages.
In this work, we extend sparse upcycling to CLIP with MoE using a single-stage training process. This approach significantly reduces training complexity and cost while achieving superior results, like shown in Figure 1. We also demonstrate its generalizability across shared or separated backbones and model scales. This establishes sparse upcycling as a practical and scalable solution for building efficient, high-performance CLIP models.
3 Sparse Upcycling with Mixture-of-Experts
Figure 2 outlines the CLIP-UP architecture and its upcycle training strategy. The left panel showcases the traditional CLIP contrastive training with expert layers replacing some MLP layers. The right panel demonstrates the innovative upcycle training approach. We’ll dive into the specifics of each in this section.
3.1 CLIP
Given $n$ pairs of image and text captions $\{(\mathbf{I}_{j},\mathbf{T}_{j})\}_{j=1}^{n}$, CLIP (Radford et al., 2021) learns image representations as $f(\mathbf{I}_{j})$ and text representation as $g(\mathbf{T}_{j})$ via contrastive learning objective within a training batch. Relevant pairs are positioned closer in the feature space, while irrelevant pairs are pushed farther apart. Given batch size of $B$ and a learned temperature parameter $\theta$, the image to text loss is
$\mathcal{L}_{I-T}=-\frac{1}{B}\sum_{j=1}^{B}\log\frac{e^{\text{sim}(f(\mathbf{I}_{j}),g(\mathbf{T}_{j}))/\theta}}{\sum_{k=1}^{B}e^{\text{sim}(f(\mathbf{I}_{j}),g(\mathbf{T}_{k}))/\theta}}$
and text to image loss is
$\mathcal{L}_{T-I}=-\frac{1}{B}\sum_{j=1}^{B}\log\frac{e^{\text{sim}(f(\mathbf{I}_{j}),g(\mathbf{T}_{j}))/\theta}}{\sum_{k=1}^{B}e^{\text{sim}(f(\mathbf{I}_{k}),g(\mathbf{T}_{j}))/\theta}}$
The contrastive loss is calculated by averaging the above two losses
$\mathcal{L}_{\text{Contrastive}}=\frac{1}{2}(L_{I-T}+L_{T-I})$
3.2 CLIP with Mixture-of-Experts upcycling
In CLIP-UP, for half of the transformer layers, we replace the dense MLPs with sparsely-gated MoE layers as an alternating dense-sparse-dense structure, similar to Figure 2 (Left). This helps the stability of feedforwarding and prevents gradient oscillation, especially during initial training.
Each MoE layer contains $E$ experts and a router. The experts are MLPs, selectively activated based on the input. The router selects the top-$K$ experts from $E$ total experts, by predicting normalized gating logits for each input token. Let $\mathbf{X}_{j}\in\mathbb{R}^{D}$ denotes the input of $j$-th token to an MoE layer, $\mathbf{G}_{j}\in\mathbb{R}^{D\times E}$ is the gating logis for $E$ experts, $\mathbf{W}_{e}\in\mathbb{R}^{D}$ is the router weight matrix for expert $e$, the output $MoE(\mathbf{X}_{j})$ is computed as follows:
$MoE(\mathbf{X}_{j})=\mathbf{X}_{j}+\sum_{e\in\text{Top-K}}g_{e,j}MLP_{e}(\mathbf{X}_{j})$
$g_{e,j}=\begin{cases}\mathbf{G}_{e,j},&\mathbf{G}_{e,j}\in\text{Top-K}({\mathbf{G}_{e,j}|1\leq e\leq E},K),\\ 0,&\text{otherwise},\end{cases}$
$\mathbf{G}_{e,j}=Softmax(\mathbf{W}_{e}^{T}\mathbf{X}_{j})$
Each expert is assigned a fixed buffer capacity (Fedus et al., 2022), meaning it can process only a limited number of tokens at a time. With predefined capacity factor $C$, and $B_{t}$ denotes the number of tokens per batch, the capacity size of each expert $B_{e}$ is computed as $B_{e}=(B_{t}/E)\times C$. This constraint enhances computational efficiency by preventing experts from being overwhelmed, while also ensuring that hardware and memory resources are managed effectively.
Tokens are assigned to experts on a ”first-come-first-serve” basis, meaning they are processed in the order they arrive (Fedus et al., 2022). This simple and effective mechanism handles input streams effectively without introducing additional complexity or overhead from token prioritization or sorting, ensuring that the system remains efficient while distributing tokens across experts.
Auxiliary loss. While the simplified token assignment reduces overhead, it presents the challenge of maintaining a balanced token distribution across all experts. Poor distribution can result in significant token dropping (Zeng et al., 2024), which can severely impact performance. Therefore, achieving balance is essential to avoid overloading or underutilizing certain experts. Overuse of some experts may lead to overfitting, while others may not receive enough data for effective training.
To mitigate this, we follow (Zoph et al., 2022) to introduce an auxiliary loss comprising load balance loss and router $z$-loss for each modality, promoting even token distribution. Specifically, the load balance loss encourages the gating mechanism to allocate tokens more uniformly over $E$ total experts. For a sequence of length $S$, the load balance loss is defined as:
$\mathcal{L}_{\text{Balance}}=\sum_{e=1}^{E}R_{e}\cdot P_{e}$
$R_{e}=\frac{E}{K\cdot S}\sum_{j=1}^{S}\mathbbm{1}(\text{Token $j$ dispatched to Expert $e$})$
$P_{e}=\frac{1}{S}\sum_{j=1}^{S}\mathbf{G}_{e,j}$
where $R_{e}$ represents the token assignment ratio to expert $e$, and $P_{e}$ denotes the average router probability of expert $e$.
The router $z$-loss is to stabilize the gating process by regularizing the output of the router logits, and it ensures the router outputs stay in a reasonable range. It is defined as:
$\mathcal{L}_{\text{Router}}=\frac{1}{S}\sum_{j=1}^{S}\left(\log\sum_{e=1}^{E}e^{\mathbf{G}_{e,j}}\right)^{2}$
The final training loss is the sum of the contrastive loss and auxiliary loss, with $\alpha$ and $\beta$ as scaling factors,
$\mathcal{L}=\mathcal{L}_{\text{Contrastive}}+\alpha\cdot\mathcal{L}_{\text{Balance}}+\beta\cdot\mathcal{L}_{\text{Router}}$
LIMOE auxiliary loss. LIMOE (Mustafa et al., 2022) proposes two new auxiliary losses, the local entropy loss and the global entropy loss, which are applied on a per-modality basis. The local entropy loss is defined as
$\mathcal{L}_{\text{Local}}=\frac{1}{S}\sum_{j=1}^{S}\left(-\sum_{e=1}^{E}\mathbf{G}_{e,j}\log\mathbf{G}_{e,j}\right)$
This encourages concentrated router weights but may reduce the diversity of expert utilization. The global entropy loss is defined as
$\mathcal{L}_{\text{Global}}=-\sum_{e=1}^{E}P_{e}\log P_{e}$
which promotes a balanced expert distribution, and enhance diversity in expert usage. To prevent a single modality from overusing too many experts, the global entropy loss is thresholded as:
$\mathcal{L}_{\text{Global}}^{\tau}=\max\{0,\tau+\mathcal{L}_{\text{Global}}\}$
where $\tau=\log(m)$, ensuring that at least $m$ experts are utilized to minimize the loss. These auxiliary losses stabilize training and improve performance, as demonstrated by (Mustafa et al., 2022).
According to the best recipe of LIMOE, applying the total training loss becomes:
$\mathcal{L}=$ $\mathcal{L}_{\text{Contrastive}}+\alpha\cdot\mathcal{L}_{\text{Balance}}+\beta\cdot\mathcal{L}_{\text{Router}}$ (11)
$+\gamma\cdot\mathcal{L}^{\text{Text}}_{\text{Local}}+\lambda\cdot\mathcal{L}^{\tau,\text{Image}}_{\text{Global}}+\delta\cdot\mathcal{L}^{\tau,\text{Text}}_{\text{Global}}$
In our experiments, we replaced the auxiliary loss with the LIMOE training recipe and tuned the hyperparameters for the best perfomrance. While the results demonstrate that the LIMOE auxiliary loss significantly enhances the performance of the shared backbone trained from scratch, its performance still lags behind other configurations and the same recipe doesn’t provide similar improvements in those settings. So we chose equation 3.2 as our final loss. More details will be illustrated in Section 4.
Table 1: Ablation study of combinations of shared vs. separated backbone and training from scratch vs. sparse upcycling. We benchmark the models on ImageNet Zero-shot classification (Accuracy@1 %); COCO and Flickr30k text-to-image (T2I) / image-to-text (I2T) retrieval (Recall@1 %).
Backbone Upcycle? Imagenet COCO Flickr30K
T2I I2T T2I I2T
Acc@1 R@1 R@1 R@1 R@1
Shared N 69.7 46.7 65.6 71.4 86.3
Shared Y 75.2 51.6 72.7 78.0 92.0
Separated N 74.5 53.1 70.6 78.3 88.2
Separated Y 76.9 52.1 71.5 80.9 92.3
3.3 Sparse Upcycling Training
The sparse upcycling process starts from a pre-trained dense CLIP model. The MoE encoders in the upcycled CLIP follow the architecture of the transformer used in (Komatsuzaki et al., 2023) including the number of layers and feature size, similarly for our separated backbone setup. A selective subset of MLP layers within the Transformer is expanded into MoE layers, with each expert initialized using the same MLP weights from the dense MLP layer of the dense CLIP model, while the router weights are randomly initialized. All other layers, such as attention and embedding layers, are copied directly from the original model. This approach ensures that only a sparse part of the model is modified, while most parameters remain unchanged, preserving the model’s core functionality. After initialization, the upcycled model undergoes additional training steps with the original hyperparameters like batch size, while slightly reducing the learning rate and weight decay to improve performance and stabilize training. The upcycle training process is illustrated in the right panel of figure 2.
Figure 3: Impact of LIMOE auxiliary loss on different training configurations. We observe adding LIMOE loss could sometimes lead to instability issue especially when the backbones are not shared, while our upcycling recipe are more robust.
4 Experiments
Datasets.
For training, we used the same paired image-text dataset for the initial dense CLIP checkpoint training, CLIP-UP, as well as the baseline dense CLIP model. The training data includes WIT-3000M (Wu et al., 2024) and DFN-5B (Fang et al., 2023). For evaluation, we used the standard ImageNet (Deng et al., 2009; Shankar et al., 2020) image classification task and COCO (Lin et al., 2014) and Flickr30K (Plummer et al., 2017) image-to-text and text-to-image retrieval tasks. The input image resolution is 224 for all of the datasets.
Setup.
We begin by training a dense CLIP model for 440k steps, which serves as the initial stage. Following this, we refine the dense model into an MoE version using the proposed upcycle training, then continuously training it for an additional 350k steps. Both stages use the AdamW optimizer with a batch size of 32k. For the dense model, we apply a learning rate of $5\times 10^{-4}$ and a weight decay of 0.2. During the upcycle training phase, the learning rate is reduced to $5\times 10^{-5}$, and the weight decay to 0.05. In the MoE model, half of the MLP layers in the Transformer are replaced with MoE layers, creating an alternating [dense, sparse] structure for both modalities. Each MoE layer includes 8 experts, with the router selecting the top 2 experts for activation. During training, we choose $\alpha=0.01$ and $\beta=0.001$ in the router auxiliary loss to ensure balanced expert distribution without dominating the primary cross-entropy objective. Hyper-parameters were fine-tuned based on preliminary experiments to improve performance and ensure training stability.
Table 2: Comparing performance of CLIP-UP and baseline models including LIMOE (Mustafa et al., 2022) across different model sizes. The sparse upcycling leverages the CLIP dense checkpoint trained for 440k steps. CLIP-UP is further trained for additional 350k steps (790k steps in total). We also train CLIP with 790k steps to have comparable training steps to CLIP-UP.
Model Steps Inference Imagenet COCO retrieval Flickr30k retrieval
classification T2I T2I I2T I2T T2I T2I I2T I2T
(k) GFLOPS Acc@1 Acc@5 R@1 R@5 R@1 R@5 R@1 R@5 R@1 R@5
B/32
OpenAI-CLIP - 14.8 63.2 88.8 30.8 55.9 51.6 75.7 - - - -
LIMOE - 22.3 67.5 - 31.0 - 45.7 - - - - -
[…]
CLIP (ours) 790 14.8 72.4 92.8 41.9 67.8 62.4 84.1 67.8 88.7 85.6 96.4
CLIP-UP 790 19.6 73.2 93.3 47.3 74.0 66.6 86.7 72.9 91.9 85.9 96.9
B/16
OpenAI-CLIP - 41.2 68.4 91.9 33.1 58.4 53.8 77.9 - - - -
LIMOE - 48.7 73.7 - 36.2 - 51.3 - - - - -
[…]
CLIP (ours) 790 41.2 76.8 95.1 44.9 70.8 66.0 86.6 74.3 92.7 88.9 98.0
CLIP-UP 790 54.3 76.9 95.1 52.1 77.6 71.5 89.2 80.9 95.6 92.3 99.2
L/14
OpenAI-CLIP - 175.5 75.3 94.5 36.1 60.8 57.7 79.1 - - - -
[…]
CLIP (ours) 790 175.5 81.6 96.6 50.2 75.2 71.4 89.9 79.3 94.9 91.7 99.0
CLIP-UP 790 231.7 81.2 96.6 53.9 79.4 73.8 92.0 82.0 96.1 92.4 99.1
Additionally, we train another dense CLIP model for a total of 790k steps to enable a fair comparison. This model is trained with the same configuration as mentioned earlier, using a learning rate of $5\times 10^{-4}$ and a weight decay of 0.2 to optimize performance.
4.1 Recipe Study
We first investigate the impact of using shared versus separated backbones and the effect of training from scratch compared to employing sparse upcycling. For a fair comparison, the shared backbone configuration uses 16 experts, matching the total number of experts (8 each for text and image) in the separated backbone setup. All experiments are conducted with the CLIP-B/16 model. All results of these models on ImageNet zero-shot classification and COCO/Flickr30k image-text retrieval tasks are illustrated in table 1.
The results show that the separated backbone combined with sparse upcycling delivers the best overall performance, achieving the highest ImageNet Top-1 accuracy of 76.9% , highest Flickr30K T2I and I2T recall@1 of 80.9% and 92.3%, and competitive results in COCO. Separated backbones generally outperform shared ones, likely because of the additional parameters dedicated to each modality. Notably, sparse upcycling provides greater improvements to shared backbones than to separated ones. Compared to shared models trained from scratch, sparse upcycling boosts ImageNet accuracy by 5.5% and COCO/Flickr retrieval by up to 7.1%, whereas the separated backbone achieves gains of 2.4% on ImageNet and up to 4.1% on retrieval tasks. Sparse upcycling proves effectiveness across both shared and separated architectures, consistently outperforming models trained from scratch. These results highlight the versatility and efficiency of CLIP-UP across different backbone configurations.
Figure 4: Performance comparison along with training EFLOPS of CLIP-UP and training from scratch on CLIP B/16.
Figure 5: CLIP-UP with MoE upcycling for only the text encoder, image encoder, or both. We observe upcycling both the image and text encoders into MoE generally helps, especially for retrieval tasks.
We further explore the impact of adding LIMOE specific auxiliary loss described in Section 3. Specifically, we set $\tau=6$ to incentivize at least 6 experts (uniformly) or more (not necessarily in a uniform way) (Mustafa et al., 2022) are used for each modality and fine-tuned the loss weight under our experiment setup. As shown in Figure 3, applying the LIMOE auxiliary loss on the shared backbone trained from scratch significantly improves both ImageNet classification and COCO retrieval metrics, which is aligned with the finding from original LIMOE paper (Mustafa et al., 2022). However, its performance remains below other configurations without applying LIMOE specific auxiliary losses. Adding LIMOE specific loss to best configuration (Separated-Upcycle) achieves a slightly better text-image retrieval metrics but falls short on ImageNet zero-shot classficaition.
In addition, we also find adding these new auxiliary losses make the model hard to train as there are much more hyper-parameters to tune. To our best effort, we cannot make LIMOE specific auxiliary loss work in all configurations, e.g. in Shared-Upcycle and Separated settings. Due to the limit of resources, we decide to use the loss in equation 3.2 as the final training objective to reduce the effort of finding best hyperparameters.
4.2 Final Model Evaluation and Baselines
According to the results from the previous section, we select the separted backbone with sparse upcycling as the default recipe and further study its effectiness by training CLIP-UP models in different model sizes, from B/32 to L/14. Table 2 compares the zero-shot image classification and retrieval performance of CLIP-UP against baseline dense CLIP model trained for an equivalent number of steps across different scales. Extending training steps for dense CLIP from 440k to 790k yields only minor gains. In contrast, CLIP-UP demonstrates substantial improvements across all metrics. These gains are consistent across model scales, particularly for image-text retrieval tasks. Remarkably, CLIP-UP B/32, using only 47% of the inference GFLOPS, outperforms the baseline dense CLIP B/16 by 2.4% in COCO T2I recall@1 and matches its I2T recall@1. Similarly, CLIP-UP B/16, requiring just 31% of the inference GFLOPS, surpasses the baseline dense CLIP L/14 by 1.9% in COCO T2I recall@1 while matching its I2T recall@1. This highlights the effectiveness of training high-performance CLIP models by leveraging pretrained dense models through sparse upcycling with MoE.
4.3 Training Efficiency
To better illustrate the effectiveness of upcycle training, we also show the training process from dense pretraining + upcycling compared to training from scratch in Figure 4. The initial dense model training provides a strong foundation for subsequent upcycling. In contrast, training from scratch without leveraging the pretrained dense checkpoint, results in much higher computational costs to reach the performance level of CLIP-UP. This difference is especially noticeable for COCO image-to-text and ImageNet classification tasks. While sparse upcycling introduces an initial quality drop on ImageNet due to model reconfiguration (Komatsuzaki et al., 2023), CLIP-UP consistently outperforms the model trained from scratch, highlighting its efficiency in reducing training resources and improving overall performance.
5 Discussion and More Ablations
In this section, we discuss important architecture and training ablations relative to the configuration used during experiment. We use the CLIP B/16 with 8 experts as the default configuration here, expert capacity factor $C=2.0$, and 6 MoE layers interspersed in the Transformer for both modalities. The hyper-parameters are same as Section 4.
5.1 MoE added to single or multiple modalities
In the CLIP-UP model, the MoE setup is applied to both the text encoder and image encoder. We also explore the effect of adding MoE layers to only one modality while keeping the other modality fully dense. The COCO retrieval and ImageNet results for these configurations are shown in Figure 5, measured across different training steps.
We observe that the initial performance of newly upcycled models tends to decline compared to the starting dense model, regardless of where the MoE setup is applied. However, all configurations recover after approximately 5k training steps. Applying MoE layers to both modalities leads to a more significant initial drop on average. When MoE layers are applied to only one modality (either image or text), the final performance remains comparable. Notably, the initial performance drop is more pronounced when MoE layers are applied to the image modality, suggesting that the model is more sensitive to changes in image representations.
5.2 Expert capacity factor
Figure 6: Effects of different expert capacity $C$.
Intuitively, the number of tokens processed by each expert plays a crucial role in determining model quality. In CLIP-UP, this is controlled by the expert capacity factor, denoted as $C$. A higher $C$ results in less token dropping, thereby reducing the initial quality drop. However, this doesn’t necessarily guarantee a higher final model quality. By default, we set the capacity factor to 2.0 for both modalities. As shown in Figure 6, increasing $C_{image}$ to 4.0 significantly boosts the ImageNet zero-shot metrics. However, this adjustment results in a noticeable drop in performance on COCO and Flickr30k retrieval tasks.
Figure 7: Visualization of image tokens dropped by the router (i.e., not being assigned to any expert due to capacity constraint) on COCO (Top) and ImageNet (Bottom).
Figure 7 visualizes the behavior of dropped image tokens under different capacity settings of COCO and Imagenet. The red squares represent the dropped image tokens. With a higher $C_{image}$, fewer image tokens are discarded. This benefits ImageNet performance since the dataset primarily consists of single-object images. Retaining more tokens allows the model to focus on the key features. In contrast, COCO images often depict complex scenes with multiple objects, but not all of which are relevant to the paired captions. Dropping less important image tokens helps the model concentrate on the most important objects, which explains the drop in COCO performance when fewer tokens are discarded. We leave the further study of capacity factor for different tasks to a future work.
6 Conclusions
In this work, we introduce CLIP-UP, a simple yet efficient training strategy for CLIP that integrates MoE with sparse upcycling. Through extensive experiments on various configurations and auxiliary losses, we show that CLIP-UP significantly reduces training costs while consistently improving retrieval and classification performance across multiple model scales. Notably, CLIP-UP outperforms larger dense models in some tasks while drastically cutting inference FLOPs, highlighting sparse upcycling as a practical and scalable approach for building efficient, high-performance CLIP models. Additionally, ablation studies provide comprehensive insights into key design decisions.
Acknowledgement
We thank Wentao Wu, Haotian Zhang, and many others for their invaluable help and feedback.
Appendix A Training Details
Below, we provide detailed training hyper-parameters and setups for dense CLIP (weights are used for sparse upcycling), sparse CLIP trained from scratch, and CLIP-UP.
A.1 Training hyper-parameters
We primarily follow (Radford et al., 2021) for hyper-parameter selection, using the WIT-3000M (Wu et al., 2024) and DFN-5B (Fang et al., 2023) training datasets. Table 3 summarizes the hyper-parameters for all experiments, including MoE-specific configurations and parameters for dense CLIP, sparse CLIP, and CLIP-UP.
Table 3: Training hyper-parameters and settings for dense CLIP used for sparse upcycling and CLIP-UP
General
Batch size 32768
Image size $224\times 224$
Text tokenizer T5 (Raffel et al., 2023), lowercase
Text maximum length 77 tokens
Optimizer AdamW ($\beta_{1}=0.9,\beta_{2}=0.98$)
LR schedule cosine decays with linear warm-up (first 2k steps)
Dropout rate 0.0
MoE
Inner structure Pre-Layer Normalization (Xiong et al., 2020)
Router type Top-2 routing
Expert capacity factor ($C$) 2.0 (both text and image)
MoE position [dense, sparse] (half of MLP layers replaced by MoE layers)
Load balance loss weight 0.01
Router-z loss weight 0.0001
Dense Model
Steps 439087 (i.e., $\sim$ 14B examples seen)
Peak learning rate (LR) $5e^{-4}$
Weight decay 0.2
CLIP-UP
Steps 351269 (i.e., $\sim$ 11B examples seen)
Peak learning rate (LR) $5e^{-5}$
Weight decay 0.05
Expert count 8 (for text and image separately)
Sparse Model
Steps 790356 (i.e., $\sim$ 25B examples seen)
Peak learning rate (LR) $5e^{-4}$
Weight decay 0.2
Expert count 8
Appendix B Ablation study
B.1 Normalize gating weights before or after routing.
To mitigate the initial quality drop observed when applying sparse upcycling, we experimented with normalizing the router output logits after routing. This ensures the remaining gating weights are normalized to sum to 1, even when some tokens are dropped due to expert capacity constraints. The intuition behind this approach is that in the dense model, each token was previously processed by a single expert MLP.
Figure 8: Model performance for gating normalization applied before or after routing
As shown in Figure 8, normalizing gating weights post-routing helps reduce the initial quality drop. However, in terms of final model performance, this approach shows improved results in image-to-text retrieval, but performs worse in text-to-image retrieval.
A possible explanation for this discrepancy is that post-routing normalization maintains the magnitude of all remaining tokens, which benefits the image encoder, as most image tokens are informative. In contrast, text encoder often deals with padding tokens, and reducing the magnitude of these tokens can enhance the text encoder’s ability to focus on meaningful content. It also aligns with the finding that the initial quality drop is the biggest when adding MoE layers into image modality only.
Appendix C Tabular results
C.1 Comparison of model architectures and impact of LIMOE auxiliary loss
All results for different model architectures, with and without LIMOE auxiliary loss, as discussed in Section 4.1.
Table 4: All results from Table 1 and Figure 3 as discussed in Section 4.1
Model Imagenet COCO Flickr30K
Accuray@1 T2I R@1 I2T R@1 T2I R@1 I2T R@1
Shared 69.7 46.7 65.6 71.4 86.3
   +LIMOE Aux. Loss 73.1 49.7 69.7 75.6 87.9
   $\Delta$ +3.4 +3.0 +4.1 +4.2 +1.6
[…]
   +LIMOE Aux. Loss 73.9 50.9 70.1 73.8 85.5
   $\Delta$1.3 -0.7 -2.6 -4.2 -6.5
[…]
   +LIMOE Aux. Loss 72.6 46.4 62.6 73.4 85.2
   $\Delta$2.0 -6.7 -8.0 -4.9 -3.0
[…]
   +LIMOE Aux. Loss 75.9 52.9 73.5 81.3 92.5
   $\Delta$1.0 +0.8 +2.0 +0.4 +0.2
C.2 Comparison of MoE added to single modality or both modalities
All results from Figure 5 to compare MoE added into different modalities.
Table 5: All results from Figure 5. MoE-text: MoE layers are added into text modality only. MoE-image: MoE layers are added into image modality only. MoE-both: MoE layers are added into both text and image modalities.
Model Steps Imagenet COCO Flickr30K
Accuracy@1 T2I R@1 I2T R@1 T2I R@1 I2T R@1
MoE-text 0 70.2 42.7 52.5 72.1 79.4
5 72.3 42.2 63.2 70.0 86.9
100 73.3 42.9 63.1 70.4 88.0
200 75.4 42.8 64.1 72.7 88.6
350 77.2 45.5 66.0 74.2 89.6
MoE-image 0 62.6 29.1 50.9 56.1 73.7
5 72.4 41.4 62.5 70.0 84.8
100 73.8 42.2 63.6 71.7 88.2
200 75.6 43.8 65.0 72.2 88.4
350 77.6 45.5 66.3 74.5 89.4
MoE-both 0 57.2 33.0 45.3 62.7 73.4
5 71.4 44.9 67.9 71.6 87.8
100 72.8 49.7 71.4 74.9 89.1
200 75.1 49.0 68.5 73.1 85.1
350 76.9 52.1 71.5 80.9 92.3
C.3 Comparison of capacity factor
All results from Figure 6 to compare the impact of capacity factor $C$.
Table 6: All results from Figure 6
Model Imagenet COCO Flickr30K
Accuracy@1 T2I R@1 I2T R@1 T2I R@1 I2T R@1
$C_{image}=2,C_{text}=2$ 76.9 52.1 71.5 80.9 92.3
$C_{image}=4,C_{text}=2$ 78.4 46.3 66.9 75.5 90.2